# JEL Classifications: C61; D15; D52; E21; G51

### v2

@@ Abstract.
Empirical evidence shows that wealthy households have substantially higher saving rates and markedly lower marginal propensity to consume (MPC) than other groups. Existing theory cannot account for this pattern unless under restrictive assumptions on returns, discounting, and preferences. This paper develops a general theory of optimal savings with preference shocks, allowing risk aversion to vary across states and over time. We show that incorporating such heterogeneity in risk attitudes fundamentally alters the asymptotic dynamics of consumption and saving. In particular, we provide an analytical characterization of the asymptotic MPCs and show that zero asymptotic MPCs, corresponding to a 100% asymptotic saving rate, arise under markedly weaker conditions than in existing theory. Strikingly, such outcomes occur whenever there is a positive probability that agents become less risk averse in the future. As a result, the vanishing MPC emerges as a generic feature rather than a knife-edge result of the optimal savings model, offering a more theoretically robust and empirically consistent account of the saving behavior of wealthy households.
JEL Classifications: C61; D15; D52; E21; G51
@@
A Theory of Saving under Risk Preference Dynamics [^1] We thank Shenghao Zhu, Nisha Peng, John Stachurski, and seminar audiences at CUEB and CUFE for helpful comments and suggestions. Qingyin Ma gratefully acknowledges financial support from the National Natural Science Foundation of China (Grant No. 72573111).
Email addresses:
qingyin.ma@cueb.edu.cn,  songxinxi@cueb.edu.cn,  alexis.akira.toda@emory.edu
Qingyin Maa,  Xinxi Songb,  Alexis Akira Todac
a,b International School of Economics and Management,
Capital University of Economics and Business
c Department of Economics, Emory University
December 28, 2025
@@ 1. Introduction
In recent decades, a large body of empirical evidence has shown that wealthy households save a substantially larger fraction of their wealth and exhibit markedly lower marginal propensity to consume (MPC) than other groups. For instance, Dynan et al. (2004) document a strong positive relationship between lifetime income and saving rates. Using Norwegian administrative data, Fagereng et al. (2025) find that households in the top 1% of the wealth distribution report saving rates far exceeding those of the median.
@@
Because household saving behavior directly drives wealth accumulation, and wealthy households hold a disproportionate share of total assets, [^2] Saez and Zucman (2016) document that the top 0.1% wealth share in the United States increased from 7% in 1979 to 22% in 2012. Similarly, Piketty et al. (2019) estimate that the top 10% wealth share in China rose from 40% in 1995 to 67% in 2015. their saving patterns have first-order implications for the long-run distribution of wealth and the mechanisms that sustain inequality (Carroll et al., 2017). From a macroeconomic perspective, even small variations in the MPC or risk-taking behavior of the wealthy can have outsized effects on aggregate saving rates, capital formation, asset prices, the transmission of fiscal and financial shocks, the equilibrium interest rate, and the design of macroprudential policy (Ampudia et al., 2024; Carroll et al., 2017; Cho, 2023; Crawley and Kuchler, 2023; Jappelli and Pistaferri, 2014; Mian et al., 2021; Patterson, 2023).
At the same time, recent research highlights that households differ not only in income, wealth, and investment opportunities, but also in their attitudes toward risk—and that these attitudes evolve over time. While the canonical incomplete market model assumes a fixed coefficient of relative risk aversion, both empirical and experimental studies reveal that risk aversion is heterogeneous across individuals and varies systematically over the life cycle and across business cycles (Andersen et al., 2008; Cohen and Einav, 2007; Falk et al., 2018; Finkelstein et al., 2013). Moreover, fluctuations in risk tolerance comove with individual financial circumstances, macroeconomic conditions, and asset market volatility (Guiso et al., 2018; Malmendier and Nagel, 2011; Sahm, 2012; Schildberg-Hörisch, 2018), suggesting that agents continuously reassess their exposure to risk in response to both individual and aggregate shocks.
These findings suggest that intertemporal consumption and saving decisions are made under dynamically evolving risk preferences, rather than under fixed utility parameters. From a theoretical standpoint, allowing for time-varying risk aversion in dynamic savings models can fundamentally reshape the asymptotic saving behavior of wealthy households, whose consumption and portfolio choices are especially sensitive to shifts in risk perceptions.
@@
Despite these empirical regularities, existing studies remain limited in explaining the persistent high saving rates and low MPCs of wealthy households. [^3] As is standard in the literature, we define the saving rate as the change in net worth relative to total income, excluding capital losses: $s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}},$ where $s_{t}$, $w_{t}$, $c_{t}$, $Y_{t}$, and $R_{t}$ denote the saving rate, wealth, consumption, nonfinancial income, and the gross rate of return on wealth, respectively. Moreover, the asymptotic saving rate refers to the saving rate in the limit as wealth $w_{t}\to\infty$. The canonical incomplete market model, as emphasized by Ma and Toda (2021), predicts that in a stationary equilibrium with constant relative risk aversion, discount factor, and return on wealth, the asymptotic saving rate is negative, contradicting the empirical evidence that the wealthy continue to accumulate assets indefinitely. Subsequent studies have introduced return risk and stochastic discounting to reconcile theory with data, yet these extensions largely preserve the assumption of constant risk aversion (Benhabib et al., 2015; Ma et al., 2020; Ma and Toda, 2021). Even in richer stochastic environments, as Ma and Toda (2021) show, matching the observed low MPCs and high saving rates among wealthy households requires quantitatively stringent conditions, and the resulting implications remain fragile.
This limitation highlights a deeper conceptual issue: the standard framework, which treats risk preferences as homogeneous and time-invariant, overlooks the endogenous interaction between risk perceptions, wealth accumulation, and saving behavior, potentially distorting predictions of saving dynamics. In this paper, we address this issue by incorporating preference shocks into a generic optimal savings model, allowing for both time-varying and state-dependent risk aversion.
@@
We then study the properties of the optimal consumption policy. Under general preference specifications, we establish its continuity and monotonicity, and derive an analytical characterization of the wealth threshold above which agents start saving. To explore the consumption and saving behavior of the wealthy households, we adopt a generalized CRRA framework. Departing from the standard formulation, we allow the level of relative risk aversion to vary with the state. This state-dependent specification aligns with the empirical evidence that individuals adjust their risk tolerance in response to changes in preferences, financial circumstances, and prevailing economic conditions, providing a more realistic foundation for studying intertemporal saving and consumption behavior.
We show that incorporating such preference heterogeneity fundamentally changes the asymptotic dynamics of saving. In particular, we provide an analytical characterization of the asymptotic MPCs and demonstrate that zero asymptotic MPCs, which correspond to a 100% asymptotic saving rate, emerge under markedly weaker conditions than in the benchmark model without preference shocks. Strikingly, when agents face a positive probability of transitioning to a lower risk-aversion state, the precautionary saving motive remains perpetually active, regardless of how much wealth has been accumulated. [^4] Indeed, the prospect of becoming less risk averse in the future induces additional precautionary saving today, as agents smooth consumption across potential preference states. In this case, consumption grows more slowly than wealth, and the asymptotic MPC converges to zero. Importantly, this result is entirely independent of stochastic discounting, return risk, or income risk, and thus represents a highly general property of asymptotic saving behavior.
The asymptotic MPC is strictly positive at some state only if the probability of transitioning to a lower risk-aversion state from that state is zero. In this case, risk attitudes evolve one-sidedly—risk aversion may remain constant or increase but cannot decline—and positive asymptotic MPCs may arise, depending on the interaction between the stochastic discount factor and the return on wealth. [^5] For the asymptotic MPCs to remain positive across all states, however, the transition matrix of risk aversion must be upper triangular, implying that agents can only become weakly more risk averse over time, which is empirically implausible. In the extreme case where agents always transition to a higher risk-aversion state (from some state), the asymptotic MPC converges to one (at that state), and the consumption function becomes nonconcave in wealth, reflecting an extreme form of precautionary behavior.
@@
Our quantitative analysis confirms that shocks to risk aversion generate economically significant and persistent responses in household consumption and saving. Importantly, both the magnitude and persistence of these responses depend on the household’s current state and on the persistence of preference shocks. In particular, the effects vary systematically with wealth and the prevailing level of risk aversion, revealing pronounced nonlinearities in the transmission of preference dynamics across the wealth distribution.
@@
Taken together, our framework provides a theoretically robust and empirically consistent account of the saving behavior of wealthy households, highlighting how risk preference heterogeneity offers a natural mechanism for the persistence of wealth accumulation. While we do not explicitly model wealth inequality, our analysis points to a novel mechanism that may be relevant for understanding wealth accumulation in richer environments.
@@
Early studies of the asymptotic saving behavior trace back to the literature on the concavity of consumption functions. Carroll and Kimball (1996) show that, under hyperbolic absolute risk aversion (HARA) preferences, the consumption function is concave in wealth, implying asymptotic linearity of consumption with respect to wealth. [^6] In fact, in a finite horizon optimal savings framework, Toda (2021a) shows that HARA preferences are also necessary for the consumption function to be concave in wealth. Carroll (2004) further emphasizes this asymptotic linearity and note that zero asymptotic MPCs can arise in a standard optimal savings model when the interest rate is negative.
@@
Building on these insights, Benhabib et al. (2015) characterize the asymptotic MPCs analytically in an optimal savings model with capital income risk, where the return to wealth follows an iid process. Ma and Toda (2021) substantially generalize this framework by allowing both the rate of return and the discount factor to be state-dependent and jointly driven by a generic Markov process. Their contribution lies in delivering a complete characterization of the asymptotic MPC in this setting and identifying conditions under which optimal consumption is asymptotically linear in wealth. In their model, zero asymptotic MPCs arise only as a knife-edge case, contingent on restrictive assumptions. In particular, their empirical analysis indicates that achieving a zero asymptotic MPC requires stringent assumptions on return risk, stochastic discounting, and key parameters such as risk aversion, limiting the empirical plausibility of their model.
@@
We extend this literature by introducing preference shocks into a generic optimal savings framework, allowing for time-varying and state-dependent risk aversion. The models in previous studies, such as Ma and Toda (2021), emerge as special cases of our framework. We show that this extension fundamentally alters the dynamics of saving and its asymptotic behavior. In contrast to Ma and Toda (2021), vanishing asymptotic MPCs are a generic feature of our framework rather than a knife-edge case. Zero asymptotic MPCs emerge naturally whenever agents face a positive probability of transitioning to a lower risk-aversion state, even without stochastic discounting and return risk. Conversely, when risk aversion can only drift upward, the consumption function becomes nonconcave in wealth and the asymptotic MPC converges to one, departing sharply from the classical concavity-based intuition of Carroll and Kimball (1996), Ma et al. (2020), and Ma and Toda (2021).
@@
Our study also builds on the theoretical literature on the optimal savings problem, a central building block of incomplete market models in modern macroeconomics (Chamberlain and Wilson, 2000; Rabault, 2002; Kuhn, 2013; Li and Stachurski, 2014; Benhabib et al., 2015; Ma et al., 2020). We extend the techniques of Li and Stachurski (2014) and Ma et al. (2020) to solve a generic optimal savings problem with preference shocks. While this literature focuses primarily on existence, uniqueness, and computability, our analysis centers on the asymptotic properties of optimal consumption, which has received little attention and has no direct counterpart in these studies.
@@
The rest of the paper is organized as follows. Section 2 formulates a generic optimal savings problem with preference shocks, and establishes existence, uniqueness, and key structural properties of the optimal policy. Section 3 systematically studies its asymptotic properties and discusses the resulting implications for saving and wealth accumulation. Section 4 presents a quantitative analysis that provides further intuition for our theoretical results. Section 5 concludes. Proofs are relegated to the Appendix.
@@ 2. Optimality results
In this section, we formulate a generic optimal savings problem with preference shocks and discuss the existence, uniqueness, and computability of the optimal consumption policy.
@@ 2.1. Problem Statement
To capture preference heterogeneity and its evolution over time, we consider a general optimal saving problem with preference shocks, where agents choose a consumption-asset path $\{(c_{t},w_{t})\}$ to solve
maximize $\mathbbm{E}_{0}\left\{\sum_{t=0}^{\infty}\left(\prod_{i=0}^{t}\beta_{i}\right)u(c_{t},Z_{t})\right\}$ (1)
subject to $w_{t+1}=R_{t+1}(w_{t}-c_{t})+Y_{t+1},$ (2)
$0\leq c_{t}\leq w_{t},$
where the initial condition $(w_{0},Z_{0})=(w,z)$ is given. Here, $u$ denotes the utility function, $\{Z_{t}\}_{t\geq 0}$ is a time-homogeneous Markov chain taking values in a finite set $\mathsf{Z}$, $\{\beta_{t}\}_{t\geq 0}$ is the discount factor process with $\beta_{0}=1$, $\{R_{t}\}_{t\geq 1}$ is the gross rate of return on wealth, and $\{Y_{t}\}_{t\geq 1}$ is non-financial income. These stochastic processes satisfy
$\beta_{t}=\beta\left(Z_{t-1},Z_{t},\varepsilon_{t}\right),\quad R_{t}=R\left(Z_{t-1},Z_{t},\varepsilon_{t}\right)\quad\text{and}\quad Y_{t}=Y\left(Z_{t-1},Z_{t},\varepsilon_{t}\right).$ (3)
@@
In this setting, $\{Z_{t}\}$ captures both preference shocks and broader idiosyncratic and aggregator factors that jointly shape agents’ intertemporal tradeoffs and consumption and saving behavior. Throughout, we define
$\mathbbm{E}_{w,z}:=\mathbbm{E}\left[\,\cdot\,\big|\,(w_{0},Z_{0})=(w,z)\right]\quad\text{and}\quad\mathbbm{E}_{z}:=\mathbbm{E}\left[\,\cdot\,\big|\,Z_{0}=z\right].$ (4)
The next period value of a random variable $X$ is typically denoted $\hat{X}$.

### v3

@@ Abstract.
Empirical evidence shows that wealthy households have substantially higher saving rates and markedly lower marginal propensity to consume (MPC) than other groups. Existing theory cannot account for this pattern unless under restrictive assumptions on returns, discounting, and preferences. This paper develops a general theory of optimal savings with preference shocks, allowing risk aversion to vary across states and over time, and shows that incorporating such heterogeneity in risk attitudes fundamentally reshapes the asymptotic dynamics of consumption and saving. In particular, zero asymptotic MPCs (100% asymptotic saving rates) arise under markedly weaker conditions than in existing theory. Strikingly, such outcomes occur whenever there is a positive probability that agents become less risk averse in the future. Therefore, the vanishing MPC emerges as a generic feature rather than a knife-edge result of the optimal savings model, offering a more theoretically robust and empirically consistent account of the saving behavior of wealthy households.
@@
A Theory of Saving under Risk Preference Dynamics*\@makefntextWe thank Émilien Gouin-Bonenfant, Mi Luo, Nisha Peng, John Stachurski, Shenghao Zhu, as well as seminar audiences at CUEB and CUFE for helpful comments and suggestions. Qingyin Ma gratefully acknowledges financial support from NSFC (No.72573111).
Qingyin Ma\@makefntextISEM, Capital University of Economics and Business. Email: qingyin.ma@cueb.edu.cn.,  Xinxi Song\@makefntextISEM, Capital University of Economics and Business. Email: songxinxi@cueb.edu.cn.,  Alexis Akira Toda§\@makefntextDepartment of Economics, Emory University. Email: alexis.akira.toda@emory.edu.
@@ 1. Introduction
In recent decades, a large body of empirical evidence has shown that wealthy households save a substantially larger fraction of their income and exhibit markedly lower marginal propensity to consume (MPC) than other groups. For instance, Dynan et al. (2004) document a strong positive relationship between lifetime income and saving rates. Using Norwegian administrative data, Fagereng et al. (2025) find that households in the top 1% of the wealth distribution report saving rates far exceeding those of the median. Moreover, Mian et al. (2025) document the emergence of a saving glut of the rich, finding that since the 1980s, the top 1% of wealthy households in US has sharply increased its saving rate from about 43% to over 54% of disposable income, while saving rates for the rest of the population have declined.
@@
Because household saving behavior directly drives wealth accumulation, and wealthy households hold a disproportionate share of total assets, [^1] Saez and Zucman (2016) document that the top 0.1% wealth share in the United States increased from 7% in 1979 to 22% in 2012. Moreover, Piketty et al. (2019) estimate that the top 10% wealth share in China rose from 40% in 1995 to 67% in 2015. their saving patterns have first-order implications for the long-run distribution of wealth and the mechanisms that sustain inequality (Benhabib and Bisin, 2018). From a macroeconomic perspective, even small variations in the MPC or risk-taking behavior of the wealthy can have outsized effects on aggregate saving rates and capital formation (Carroll et al., 2017; Lee, 2023), the equilibrium interest rates and asset prices (Mian et al., 2013, 2021; Patterson, 2023), the transmission of fiscal and financial shocks (Cho, 2023; Ampudia et al., 2024; Auclert et al., 2025), and the design of macroprudential policy (Jappelli and Pistaferri, 2014; Crawley and Kuchler, 2023).
@@
Despite these empirical regularities, existing studies remain limited in explaining the persistent high saving rates and low MPCs of wealthy households. [^2] As is standard in the literature, we define the saving rate as the change in net worth relative to total income, excluding capital losses: $s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}},$ where $s_{t}$, $w_{t}$, $c_{t}$, $Y_{t}$, and $R_{t}$ denote the saving rate, wealth, consumption, nonfinancial income, and the gross rate of return on wealth, respectively. The asymptotic saving rate refers to the saving rate in the limit as $w_{t}\to\infty$. The canonical incomplete market model, as emphasized by Ma and Toda (2021), predicts that in a stationary equilibrium with constant relative risk aversion, discount factor, and return on wealth, the asymptotic saving rate is negative, contradicting the empirical evidence that the wealthy continue to accumulate assets indefinitely. Subsequent studies have introduced return risk and stochastic discounting to reconcile theory with data (Benhabib et al., 2015; Ma et al., 2020; Ma and Toda, 2021). However, even in richer stochastic environments, as Ma and Toda (2021) show, matching the observed low MPCs and high saving rates among wealthy households requires quantitatively stringent conditions, and the resulting implications remain fragile.
A key limitation of this literature lies in treating risk preferences as homogeneous and time-invariant. Recent research highlights that households differ not only in income, wealth, and investment opportunities, but also in their attitudes toward risk—and that these attitudes evolve over time. While the canonical savings model assumes a fixed coefficient of relative risk aversion, a growing body of empirical and experimental studies reveal that risk aversion is heterogeneous across individuals and varies systematically over the life cycle and across business cycles (Andersen et al., 2008; Cohen and Einav, 2007; Falk et al., 2018; Finkelstein et al., 2013). Moreover, fluctuations in risk tolerance comove with individual financial circumstances, macroeconomic conditions, and asset market volatility (Guiso et al., 2018; Malmendier and Nagel, 2011; Sahm, 2012; Schildberg-Hörisch, 2018), suggesting that agents continuously reassess their exposure to risk in response to both individual and aggregate shocks.
These observations motivate a reconsideration of the intertemporal saving behavior under dynamically evolving risk preferences. By abstracting from fluctuations in risk attitudes, existing studies overlook the endogenous interaction between risk perceptions, wealth accumulation, and saving behavior, potentially distorting predictions of long-run saving dynamics. [^3] From a theoretical perspective, the assumption of constant risk aversion plays a critical role in generating the negative asymptotic saving rates implied by standard models. With fixed preferences, precautionary motives weaken as wealth grows, causing consumption to eventually outpace income and leading saving rates to turn negative at high wealth levels. Allowing risk aversion to vary over time and across states, by contrast, can profoundly alter consumption and saving incentives, particularly for wealthy households, whose consumption and portfolio choices are especially sensitive to shifts in risk perceptions.
In this paper, we develop a general theory of optimal savings with preference shocks, allowing risk aversion to be both time-varying and state-dependent, and show that this generalization fundamentally reshapes the asymptotic dynamics of consumption and saving, offering a parsimonious and robust explanation for the persistently high saving rates and low MPCs observed among wealthy households.
@@
We then study the properties of the optimal consumption policy. Under general specifications, we establish its continuity and monotonicity, and analytically characterize the wealth threshold above which saving occurs. To explore the consumption and saving behavior of the wealthy households, we adopt a generalized CRRA framework. Departing from the standard formulation, we allow the level of risk aversion to vary with the state. This state-dependent specification aligns with the empirical evidence that individuals adjust their risk tolerance in response to changes in preferences, financial circumstances, and prevailing economic conditions, providing a more realistic foundation for studying intertemporal consumption and saving behavior.
We show that incorporating such preference heterogeneity fundamentally changes the asymptotic dynamics of saving. In particular, we provide an analytical characterization of the asymptotic MPCs and demonstrate that zero asymptotic MPCs, which correspond to a 100% asymptotic saving rate, emerge under markedly weaker conditions than in the benchmark model without preference shocks. Strikingly, whenever agents face a positive probability of transitioning to a lower risk aversion state, the precautionary saving motive remains perpetually active, regardless of how much wealth has been accumulated. [^4] Indeed, the prospect of becoming less risk averse in the future induces additional precautionary saving today, as agents smooth consumption across potential preference states. In this case, consumption grows more slowly than wealth, and the asymptotic MPC converges to zero. Importantly, this outcome is entirely independent of stochastic discounting, return risk, or income risk, and thus represents a highly general property of asymptotic saving behavior.
The asymptotic MPC is strictly positive at some state only if the probability of transitioning to a lower risk aversion state from that state is zero. In this case, risk attitudes evolve one-sidedly—risk aversion may remain constant or increase but cannot decline—and positive asymptotic MPCs may arise, depending on the interaction between the stochastic discount factor and the return on wealth. [^5] For the asymptotic MPCs to remain positive across all states, however, the transition matrix of risk aversion must be upper triangular, implying that agents can only become weakly more risk averse over time, which is empirically implausible. In the extreme case where agents always transition to a higher risk aversion state (from some state), the asymptotic MPC converges to one (at that state), and the consumption function becomes nonconcave in wealth, reflecting an extreme form of precautionary behavior.
@@
Our quantitative analysis confirms that shocks to risk aversion generate economically significant and persistent responses in consumption and saving. Importantly, both the magnitude and persistence of these responses depend on households’ current state and on the persistence of preference shocks. In particular, the effects vary systematically with wealth and the prevailing level of risk aversion, revealing pronounced nonlinearities in the transmission of preference shocks across the wealth distribution.
@@
Taken together, our framework provides a theoretically robust and empirically consistent account of the saving behavior of wealthy households, highlighting how risk preference heterogeneity offers a natural mechanism for the persistence of wealth accumulation. While we do not explicitly model wealth inequality, our analysis points to a novel mechanism that may be relevant for understanding wealth distribution in richer environments.
Related Literature
@@
Early studies of the asymptotic saving behavior trace back to the literature on the concavity of consumption functions. Carroll and Kimball (1996) show that, under hyperbolic absolute risk aversion (HARA) preferences, the consumption function is concave in wealth, implying asymptotic linearity of consumption with respect to wealth. Carroll (2004) further emphasizes this asymptotic linearity and note that zero asymptotic MPCs can arise in a standard optimal savings model when the interest rate is negative.
@@
Building on these insights, Benhabib et al. (2015) characterize the asymptotic MPCs analytically in an optimal savings model with capital income risk, where the return to wealth follows an iid process. Ma and Toda (2021) generalize this framework by allowing both the rate of return and the discount factor to be state-dependent and jointly driven by a generic Markov process. Their contribution lies in delivering a complete characterization of the asymptotic MPC in this setting and identifying conditions under which optimal consumption is asymptotically linear in wealth. In their model, zero asymptotic MPCs arise only as a knife-edge case, contingent on restrictive assumptions. Their empirical analysis indicates that achieving a zero asymptotic MPC requires stringent assumptions on return risk, stochastic discounting, and key parameters such as risk aversion, limiting the empirical plausibility of their model.
@@
We extend this literature by incorporating preference shocks into a generic optimal savings framework, allowing for time-varying and state-dependent risk aversion. The models in previous studies, such as Ma and Toda (2021), emerge as special cases. In sharp contrast to these studies, the vanishing asymptotic MPCs are a generic feature of our framework rather than a knife-edge case. Furthermore, when risk aversion can only drift upward, consumption becomes nonconcave in wealth and the asymptotic MPC converges to one, departing sharply from the classical concavity-based intuition of Carroll and Kimball (1996), Ma et al. (2020), and Ma and Toda (2021).
@@
Our study also builds on the theoretical literature on the optimal savings problem, a cornerstone of modern macroeconomics (Chamberlain and Wilson, 2000; Kuhn, 2013; Li and Stachurski, 2014; Benhabib et al., 2015; Ma et al., 2020). We extend the method of Li and Stachurski (2014) and Ma et al. (2020) to solve a generic savings problem with preference shocks. While this literature focuses primarily on existence, uniqueness, and computability, our analysis centers on the asymptotic properties of consumption and saving, which has no direct counterpart in these studies.
@@
The rest of the paper is organized as follows. Section 2 formulates a generic optimal savings problem with preference shocks, and establishes existence, uniqueness, and key properties of the optimal policy. Section 3 systematically studies its asymptotic properties and discusses the resulting implications for saving and wealth accumulation. Section 4 presents a quantitative analysis that provides further intuition for our theory. Proofs and detailed discussion of the related algorithms are relegated to the Appendix.
@@ 2. Optimality results
To capture preference heterogeneity and its evolution over time, we study a general optimal savings problem with preference shocks.
@@ 2.1. Problem Statement
Agents choose a consumption-asset path $\{(c_{t},w_{t})\}$ to
(1) maximize $\mathbbm{E}_{0}\left\{\sum_{t=0}^{\infty}\left(\prod_{i=0}^{t}\beta_{i}\right)u(c_{t},Z_{t})\right\}$
(2) subject to $w_{t+1}=R_{t+1}(w_{t}-c_{t})+Y_{t+1}\quad\text{and}\quad 0\leq c_{t}\leq w_{t}.$
Here, the initial condition $(w_{0},Z_{0})=(w,z)$ is given, $u$ denotes the utility function, $\{Z_{t}\}_{t\geq 0}$ is a time-homogeneous Markov chain taking values in a finite set $\mathsf{Z}$, $\{\beta_{t}\}_{t\geq 0}$ is the discount factor process with $\beta_{0}=1$, $\{R_{t}\}_{t\geq 1}$ is the gross rate of return on wealth, and $\{Y_{t}\}_{t\geq 1}$ is non-financial income. These stochastic processes satisfy
(3) $\beta_{t}=\beta\left(Z_{t-1},Z_{t},\varepsilon_{t}\right),\quad R_{t}=R\left(Z_{t-1},Z_{t},\varepsilon_{t}\right)\quad\text{and}\quad Y_{t}=Y\left(Z_{t-1},Z_{t},\varepsilon_{t}\right),$
@@
In this setting, $\{Z_{t}\}$ captures both preference shocks and broader idiosyncratic and aggregator factors that jointly shape agents’ intertemporal tradeoffs and consumption and saving behavior. To simplify notation, we denote the partial derivative of utility with respect to consumption by
$u^{\prime}(c,z):=\partial u(c,z)\big/\partial c.$
In addition, we denote $\hat{X}$ as the next period value of a random variable $X$, and
(4) $\mathbbm{E}_{w,z}:=\mathbbm{E}\left[\,\cdot\,\big|\,(w_{0},Z_{0})=(w,z)\right]\quad\text{and}\quad\mathbbm{E}_{z}:=\mathbbm{E}\left[\,\cdot\,\big|\,Z_{0}=z\right].$